Answers?
For the yi PROBLEMIV 1 Points pes e in ctheoof Goods Sld schedule tor Part 1: Review the t Part 2: Caloulate the lower of cost or or market for 2018 using the direct me Exhibit A: Ending Inventory Count documents in the Exhibits belowng ora, Incs 2018 financial statements Seckorss 2018 ieventory to value ending inventory at the lower of cost OFFICE MEMO INTERO To: Accounting Manager From: Controller Re: 2017 Ending Inventory Count 0t 147 s of the annual chysical inventory as of December 31, 2017 were 20,000 as of units at $7.50 per unit Exhibit B: Inventory Footnote Disclosure NOTE 4: INVENTORIES Inventories consist primarily of one product whchw Inventories are valued at the lower of cost or market. We determine last-in, first-out (LIFO) method. $9.20 per unit and the net realizable value was $1000 per unit. Our normal profit margin is S1.20 per unit. purchase from various suppliers cost on the basis of the At December 31, the replacement cost of our inventory was Exhibit C: Purchase Journal kora, Inc. - Purchase Journal Freight Purchase Accounts Units Purchases In Discounts Payable 15,000 115, 000 5,75013,450)117,300 18, 000 150, 000 7,500 4, 500) 153,000 12,000 120,000 6,000 (3,600) 127,500 25, 000 175,000 8,750 (5,250) 178, 500 Date March 31 June 30 September 30 December 31 Total 000 560,00026,000 116t B001 571,200 一 Exhibit D: Sales Journal Seckora, Inc. -Sales Journal Freight Sales Accounts Date Sales Out Discounts Receivable 33, 750) 75,000 1.125,000 5t5 Total
Answer #1
1 Cost of goods sold schedule:
Uder LIFO method, We assume that goods purchased last are sold first
Here, 75000 units are sold in the following order:
Date Units Total
cost
Dec 31. 25000 178500
Sep 30. 12000 127500
Jue 30. 18000 153000
Mar 31. 15000 117300
Jan 1. 5000 37500
(Beg. Inv) (5000*7.50)
75000 613800
Cost of goods sold=$ 613800
2 Cost of inventory will be the balance units remaining in the beginning inventory
Cost=15000*7.5=$ 112500
Market value:
Replacement cost of $ 9.20 shall be taken as market value subject to ceilig ad floor limit
Ceiling limit=Net realizable value=$ 10 per unit
Floor limit=Ceiling limit-normal profit margin=10-1.20=$ 8.80 per unit
Replacement cost shall come within ceiling and floor limit
$ 9.20 comes within $ 10 and $ 8.80
Hence,market value=Replaement cost=$ 9.20
Total value of inventory using market value=15000*9.20=$ 138000
Lower of cost or market= Lower of $ 112500 and $ 138000= $ 112500 (at cost)
3 No adjusting etry required to make since the inventory are valued at cost